OpenStudy (anonymous):

A company builds computers. It costs $7,600 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? C(x) = 460x − 3000 C(x) = 3000x − 460 C(x) = 460x + 3000 C(x) = 3000x + 460

OpenStudy (anonymous):

that one is hard............

OpenStudy (anonymous):

i know, i dont know what to do...

OpenStudy (anonymous):

what if i show u how to do it with another problem

OpenStudy (anonymous):

okay

OpenStudy (ganpat):

say a is basic cost and b is cost for computer.. so a + 10b = 7600 and a + 20b = 12200 2a + 20b = 15200 a + 20 b = 12200 subtract a = 3000 and b = 460.. so C(x) = a + xb = 3000 + 460x

OpenStudy (anonymous):

oh good job !

OpenStudy (anonymous):

thanks !

OpenStudy (ganpat):

w/c.. thnx !! :)